Find the exact values of cos theta and sin theta for the angle measure.
-3 pi/4 radians

Answer by stanbon(75887)

Find the exact values of cos theta and sin theta for the angle measure.
-3 pi/4 radians
----
cos(-3/4 pi) = -sqrt(2)/2
sin(-3/4 pi) = -sqrt(2)/2
============================
Cheers,
Stan H.
============================